OpenStudy (anonymous):

Find the indefinite integral of: [1 + 4cotx]/[4 - cotx] dx I would like a hint please

OpenStudy (anonymous):

I supposed ∫ csc^4 x cot^6 x dx = - ∫ 1/sin^2 x cot^6 x d(cot x) = - ∫ (sin^2 x + cos^2 x)/sin^2 x * cot^6 x d(cot x) = - ∫ (1 + cot^2 x) cot^6 x d(cot x) = - ∫ ( cot^6 x + cot^8 x) d(cot x) = -(1/7)cot^7 (x) - (1/9)cot^8 (x) + C Or -(1/7)(cot (x))^7 - (1/9)(cot (x))^9 + C

sam (.sam.):

Trig sub, \[\text{Let} ~u=\tan \left(\frac{x}{2}\right)\] ,\[and~~du=\frac{1}{2} \sec ^2\left(\frac{x}{2}\right) d x\]

sam (.sam.):

You'll get \[\huge \int\limits \frac{2 \left(1-\frac{2 \left(u^2-1\right)}{u}\right)}{\left(u^2+1\right) \left(\frac{u^2-1}{2 u}+4\right)} \, du\] Simplifying \[\text{}=\int\limits \frac{-8 u^2+4 u+8}{u^4+8 u^3+8 u-1} \, du\] Seaparate using partial fractions, \[\text{}=\int\limits \left(-\frac{2 u}{u^2+1}+\frac{1}{u-\sqrt{17}+4}+\frac{1}{u+\sqrt{17}+4}\right) \, du\] Solve

sam (.sam.):

You can check your final answer with \[\ln \left(\tan \left(\frac{x}{2}\right)-\sqrt{17}+4\right)+\ln \left(\tan \left(\frac{x}{2}\right)+\sqrt{17}+4\right)-\ln \left(\sec ^2\left(\frac{x}{2}\right)\right)+c\]

OpenStudy (anonymous):

How were you able to use tanx? All I see is cotx...